Please help me solve for k
6(k-1)-(5k-9)= -6
I have tried solving it as follow:
6k-6-5k-9= -6
-5k -5k
k-15= -6
+15 +15
k=9

Answer by rothauserc(4718)

6(k-1)-(5k-9)= -6
6k-6-5k+9 = -6
k+3 = -6
k = -9
check the answer
6(-9-1)-(5*-9-9) = -6
-60 -(-54) = -6
-60+54 = -6
-6 = -6
answer checks :-)